OpenStudy (anonymous):

Solve the following system. x 2 + y 2 = 25 2x + y = 10 The solution set {(________,__________,_________,________)}

OpenStudy (anonymous):

i need 4 answers!!!

OpenStudy (anonymous):

but how would you do that i only got 2

OpenStudy (welshfella):

use the substitution method from the second equation y = 10 - 2x Now replace y in the first equation by 10-2x to get an equation in x only solve this for x ( there will be 2 values of x)

OpenStudy (anonymous):

this is what i got 2x+y=10;2x+2y=25 Step: Solve 2x+y=10 for y: 2x+y+−2x=10+−2x(Add -2x to both sides) y=−2x+10 Step: Substitute (−2x+10) for y in 2x+2y=25: 2x+2y=25 2x+2(−2x+10)=25 −2x+20=25(Simplify both sides of the equation) −2x+20+−20=25+−20(Add -20 to both sides) −2x=5 −2x−2=5−2(Divide both sides by -2) x=−52 Step: Substitute (−52) for x in y=−2x+10: y=−2x+10 y=−2(−52)+10 y=15(Simplify both sides of the equation) Answer: y=15 and x=−52

OpenStudy (welshfella):

you can then find the corresponding values of y by substituting back in equation 2

OpenStudy (anonymous):

mm

OpenStudy (anonymous):

so can you solve this for me please and then explain after

OpenStudy (anonymous):

Im confused

OpenStudy (welshfella):

do you follow my method?

OpenStudy (anonymous):

Not really but please can you do it and explain?

OpenStudy (welshfella):

2x + y = 10 subtract 2x from both sides y = 10 - 2x OK with that?

OpenStudy (anonymous):

yes i am

OpenStudy (welshfella):

right x^2 + y^2 = 25 replace the y in the above equation by 10 - 2x:- x^2 + (10- 2x)^2 = 25 x^2 + (10 - 2x)(10 - 2x) = 25 expanding the brackets:- x^2 + 100 -20x-20x + 4x^2 = 25 adding like terms 5x^2 - 40x + 100 = 25 5x^2 - 40x + 75 = 0

OpenStudy (anonymous):

alright

OpenStudy (welshfella):

so now we need to solve this quadratic you can make it easier by dividing each term by 5:- giving x^2 -8x + 15 = 0

OpenStudy (welshfella):

can you solve this? you need to factor first you need 2 numbers whose product is 15 and when added = -8

OpenStudy (anonymous):

just do it foe me sir i have limited time

OpenStudy (welshfella):

(x - 3)(x - 5) = 0 x = 3 or 5 now plug x=3 into the second equation 2x + y = 10 and find y then do the same for x = 5 and you have your 4 solutions

OpenStudy (anonymous):

i dont have a calculator

OpenStudy (welshfella):

you dont need one to these

OpenStudy (anonymous):

i got 4,8,10 and 12 is that correct

OpenStudy (welshfella):

2(3) + y = 10 6 + y = 10 y = ?

OpenStudy (anonymous):

if not what are they

OpenStudy (welshfella):

if already got 2 values of x :- 3 and 5 now we need values of y

OpenStudy (welshfella):

6 + y = 10 y = ? i'm sure you can do that

OpenStudy (anonymous):

its 4

OpenStudy (welshfella):

right now plug in x = 5 to find the second value of y

OpenStudy (anonymous):

0?

OpenStudy (welshfella):

yes

OpenStudy (welshfella):

he has the answers now

OpenStudy (anonymous):

so its 3,5,4,0?

OpenStudy (anonymous):

lol my fault

OpenStudy (welshfella):

yes but pair them up ( x,y) (5,0) and (3,4)

OpenStudy (anonymous):

but i have 4 answer boxes?

OpenStudy (welshfella):

well enter then as 5 0 3 4

OpenStudy (anonymous):

look at the flick of the wrist

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np

OpenStudy (welshfella):

yw